40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 161 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  9213   Mon Oct 7 13:55:26 2013 JenneUpdateLSCArms locked in IR for many hours

Someone left the arms aligned, and the LSC engaged, so the arms have been locked almost continuously for several days hours.  The trend below is for 4 days hours.  What is most impressive to me is that we don't see a big degredation in the transmitted power over this time.

EDIT: Okay, I got excited without paying attention to units.  It was only several hours, which is not too unusual.  Although the lack of transmission degredation is still unusual.  However, this may be due to improved oplevs?  I'm not sure why, but we're not seeing (at least in this plot) the degredation to ~0.7 after an hour or so.

BothArmsLocked3Days_7Oct2013.png

  9212   Mon Oct 7 10:51:18 2013 SteveUpdateSUSReplaced the laser for Optical Lever of ETMY

 Just plot.

RA: I'm not sure how to interperet this; I think that the SUM channel is divided by the SUM so that this is supposed to be RIN, but not sure. Can someone please take a look into the SUS model and then explain in the elog what the SUM normalization algorithm is?

Attachment 1: ETMXoplevETMY.png
ETMXoplevETMY.png
Attachment 2: oplevSettings.png
oplevSettings.png
  9211   Sun Oct 6 23:50:14 2013 ranaConfigurationSUSMC filters adjusted
  1. Found the low pass filters OFF in a couple of the MC SUS damping loops. This allows injection of OSEM noise all the way up to ~100 Hz. I deleted unused ones and turned on the 'Cheby' for all SUS DOFs: cheby1("LowPass",2,0.1,3)cheby1("LowPass",6,1,12)gain(1.13501)
  2. Tuned the Bounce/Roll filters to catch the bounce and roll modes for all 3 MC SUS (based on the Mech Res Wiki page). These are now engaged on ALL MC SUS DOFs. They have been deleted from the ASCPIT/YAW filter banks and moved into the WFS screen into the MC1,2,3 filter banks there to be more in line with our knowledge of multi loop instability from notches in individual loops:ellip("BandStop",4,1,40,15.9,17.2)ellip("BandStop",4,1,40,23.5,24.7)gain(1.25893)
  9210   Sun Oct 6 23:43:07 2013 ranaUpdateIOOWFS debugging

Quote:

Tried a bunch of stuff, but eventually just turned off the TRANS_QPD loops and loops are stable. Needs more debugging.

  1. Modified the on/off scripts so that the Integrators are no longer toggled. No reason to turn them off since we are clearing the filter bank histories.
  2. With QPD feedback OFF, I have lowered the overall gain by 15x so that its just drift control.
  3. Deleted unused / bad filters from the main filter banks.
  4. Gautam is going to debug the QPD with a red laser pointer and then elog.
  5. Jamie is checking out the MC Coil dewhitening logic to see if that's in a funny state.

 Back around June 18, Jamie was debugging some Guardian code here to replace our MC autolocker. Afterwards our MC WFS stopped working. We never figured out what went wrong, but at the time we turned off the feedback from the MC trans QPD and it stabilized the response at DC.

Today, I noticed that the trans QPD feedback is on.  Did anyone do this on purpose?

Its problem causing behavior is slow, but you can catch it if you wait. With the nominal WFS gain of 0.4 the control signal ramps up monotonically at a rate of ~100 counts/minute. Depending upon the static alignment of the MC, this could let it take 10 minutes or a few hours before it rails the MC SUS actuators and breaks the lock. Very sneaky. Don't turn this loop back on without making sure its working and not breaking. I would trend it for you, but the SLOW channels associated with the TRANS QPD servo are not trended --- does anyone know how to get them in the channel list?

  9209   Sun Oct 6 22:52:09 2013 Jenne, RanaUpdateIOOinput beam to PMC aligned again

pmcr.pngafter

I wonder what's drifting between the laser and the PMC? And why is it getting worse lately?

  9208   Sun Oct 6 22:27:35 2013 ranaFrogsElectronicsMC3 LL sensor cable was loose

I noticed that the MC3 LL sensor was apparently dead according to its suspension screen. Since it was only the fast ADC channel and not the SLOW PDmon, I could tell that it was just in the ADC cabling. I pushed in a few of the MC3 sensor cables on the front and back of the PD whitening board and it came back OK. According to this trend of the past 40 days and 40 nights, it started slipping on this past Wednesday morning.

Was anyone walking near MC2 or the suspension electronics racks before noon on Wednesday (Oct. 2nd)?

Attachment 1: MC3_LL.png
MC3_LL.png
  9207   Sun Oct 6 20:55:08 2013 ranaSummaryASCMC WFS Limits set based on 40 days of trends

MC3 watchdog gets tripped sometimes when lock is lost. I noticed that there were no limits set in the MC WFS drive. The attached plot shows that over 40 days, the OUT16 channels from the WFS don't exceed 1000 counts. So I've set the limit to be 2000 in all 6 of the MC ASCPIT/YAW filter banks. Please don't turn them off.

OUT16 is really not the right way to measure this, but for some reason, we don't have any DQ channels from the MC WFS screen ??? So we're not able to measure the trend of the high frequency drive signal.

So I added the WFS(1,2)_I_(PIT,YAW)_OUT_DQ and WFS(1,2)_(PIT,YAW)_OUT_DQ channels to the c1ioo.mdl at 2048 Hz. I used Jamie's excellent 'rtcds' utility to build and install:

1) after making the edits to c1ioo.mdl I saved the file/

2) sshing to c1ioo

3) rtcds stop c1ioo

4) rtcds make c1ioo

5) rtcds install c1ioo

6) rtcds start c1ioo

7) telnet fb 8087

8) daqd> shutdown

That seemed to do it OK.

Unfortunately, all of the instructions that we have in the Wiki for adding channels and model building are misleading and don't mention any of this. There are a few different methods listed which all instruct us to do the whole make and make install business in a bunch of non existent directories.

Attachment 1: mcwrfs_trend.png
mcwrfs_trend.png
  9206   Sun Oct 6 18:37:43 2013 ranaUpdateSUSReplaced the laser for Optical Lever of ETMY

I centered the ETMY OL today and found that the UGF was around 3-4x too LOW after the laser swap and re-alignment. That's why the Y arm has been shaking so much today.

NO more OL work without loop measurements and noise measurements.

web-burnt-toast.jpg

  9205   Sun Oct 6 17:05:49 2013 ranaSummaryIOOMC ASC problems

MC unlocked over the weekend and also got severely mis-aligned. It all started around midnight on Saturday.

At first I thought that this was due to the MCS CPU meter being railed at 60 us, so I deleted a bunch of filters in MC1,2,3 that are unused and left over from Den's quantization noise investigations. This reduced the CPU load somewhat, but didn't make any real improvements. Turning on the ASC filter banks in the MC SUS still mis-aligned the MC.

With the MC WFS and MC ASS turned off, there is still some digital junk coming in and misaligning things. Plot attached.

Similar stuff coming in on ITMX, but not ITMY.

Tried restarting various FEs, but there was no effect. Also tried rebooting c1lsc, c1ioo, & c1sus. Finally did 'shutdown -r now' on all 5 computers on the CDS overview screen and simultaneously (almost) pressed the reset button on the RFM switch above the old c1pem crate. Everything came back OK except for c1oaf (I had to manually button his BURT button) and now the ASC inputs on all the SUS are zero when they should be and MC is well locked and aligned.

Rob and I used to do this trick when he thought that a cosmic ray had corrupted a bit in the RFM network.

Attachment 1: mcasc.pdf
mcasc.pdf
  9204   Fri Oct 4 20:25:12 2013 MasayukiUpdateGreen LockingALS down script

I wrote the down script for ALS. This script is  (script)/ALS/ALSdown.py When this script is running, it watches the feedback signal of the ALS control loop so as to shut down the servo immediately when  the suspension is kicked. 

When  the value of  C1:ALS-X(Y)ARM_OUT  becomes larger than the threshold (right now it is 4500 counts), it changes the servo gain to 0, turns off all filters except for FM5 (the filter for phase compensation), resets the history of the phase tracker of each arm and prints the time on window when the suspension kicked

I put the switch on the C1ALS screen, and if you push this switch the window will open (like when you turn on the c1ass script) and the script start to run. For stopping this script, you have to close that window or press Ctrl + C on that window. This is little bit inconvenient, but we will make autolocker script for ALS and this downscript will be included that script soon. So I think it is enough to protect the suspensions right now.

Attachment 1: Screenshot-Untitled_Window.png
Screenshot-Untitled_Window.png
  9203   Fri Oct 4 14:36:44 2013 ranaUpdateSUSReplaced the laser for Optical Lever of ETMY

  That's good, but please no more Oplev work. We want to do all of it at once and to make no more changes until we have all the parts (e.g. dumps and correct lenses) in hand and then talk over what the new design will be. I don't want to tune the beam size and loop shape every week.

  9202   Fri Oct 4 12:33:27 2013 MasayukiUpdateSUSReplaced the laser for Optical Lever of ETMY

[Steve, Masayuki]

We replaced the laser for optical lever of ETMY. And also we aligned the path so that beam spot hits the center for each optics. I attached the spectrum of the SUS-ETMY_OPLEV_SUM, the red curve is with old laser, and blue curve is with the new laser. We also measured without laser so as to measure the QPD dark noise (green curve). The change is significant, and seems much closer to other oplev spectrum.(Brown curve is the oplev spectrum of ITMY)

The new laser is:

Manufacture name: JDSU, Model number: 1103P, Serial number: PA892324

The injection power is 3.7 mW and the out coming power is 197 uW (measured just before the QPD). The output value of the SUS-ETMY_OPLEV_SUM is about 8500.

First we measure 2 spectrum ( including the dark noise). After that we replace the laser and align the optical lever path. We changed the post of one of the mirror (just before the QPD) because the hight was too low. Inside of the chamber is darker than before - actually we had scattering light inside the chamber before. We dumped the reflected light from QPD. And then we measured the spectrum of the oplev output. I also aligned oplev of ETMY after restoring the YARM configuration using IFO configure screen.

We don't know actually what was the problem, laser quality or the scattering light or some clipping. But the oplev seems to be better.

Steve:  Atm2 shows increased gains correction later for UGF elog 9206

Attachment 1: OPLEV_SUM.pdf
OPLEV_SUM.pdf
Attachment 2: ETMYoplev.png
ETMYoplev.png
  9201   Fri Oct 4 02:08:32 2013 MasayukiUpdateGreen LockingFPMI with ALS arm stabilization

[Manasa, Masayuki]

We locked MICH with 2 arms stabilized by ALS control.

Measurement

We measured the power spectrum of the LSC-MICH_IN1  at each step so as to know the in-loop noise of MICH. And also we measured the OLTF of MICH loop and the error signal with BS excited at 580 Hz and MICH notch filter at same frequency enabled to obtain the MICH calibration factor.

1. We locked  MICH using the AS55Q error signal and fedback to BS actuator. (Red curve)

2. We locked  MICH and locked both the arms using POX11 and POY11 error signals and fedback to ETMs actuators.(Blue curve)

3. We stabilized both the arms using ALS. We use the ALS error signals and fedback to ETMs actuators. And then we locked  MICH.(Magenta curve)

Attachment

The green and brown curve are the ALS in-loop noise, which is the _PHASE_OUT_Hz calibrated error signals. So for these two curves the unit of vertical axis is Hz/rHz. The other curves are the MICH in-loop noises and these are not calibrated. So for these curves the unit of vertical axis is counts/rHz.

Discussion

The UGF of MICH loop is 10 Hz with phase margin of 45 degrees (measured today). The FPMI noise with ALS stabilized arms is much larger than the FPMI with IR PDH locked arms above 30 Hz. That is because the ALS arm stability is not as good as the stability of PDH locked arms. We have to analyze and verify the calibrated numbers for FPMI + ALS with model.

Attachment 1: FPMI_ALS.pdf
FPMI_ALS.pdf
  9200   Fri Oct 4 01:09:33 2013 ManasaUpdateGreen LockingFrequency counter for ALS

I checked the BK precision 1856D manual. I found that although this frequency counter can measure upto 3.5GHz, it has 2 separate input channels to measure two range of frequencies.

One input to measure between 0.1Hz to 100 MHz and the other to measure between 80MHz to 3.5GHz. Our beat frequency desirable range is <100MHz for stable ALS. Also, the beat PD response falls off beyond ~150MHz . Should we be happy with this frequency counter and use it in the 0.1Hz-100MHz range or look for one with a better measuring range?

P.S. Right now we are using the spectrum analyzer in the control room set to frequency range from 10MHz - 140 MHz for beat note search. 

  9199   Thu Oct 3 22:14:06 2013 MasayukiUpdateGreen LockingBoth arms ALS stabilized for IR resonance

[Manasa, Masayuki]

We succeeded in stabilizing both the arms using ALS and get IR to resonate at the sametime.

Measurement

At each step we measured the _PHASE_OUT_Hz calibrated error signals for  Y in this configuration so as to get the in-loop noise of ALS control of YARM

1. we stabilized YARM off IR resonance by using ALS, misaligned ETMX, closed XARM green shutter. That means no IR flashing and no green in XARM.

2. we aligned the ETMX with XARM green shutter closed.

3. we opened the green shutter and locked the green laser with PDH to the XARM.

4. we stabilized the XARM using ALS and off resonance for IR.

5.We brought the XARM to IR resonance with YARM stabilized off IR resonance.

6. we brought the YARM to IR resonance

Beat frequencies when both the arms were stabilized and had IR resonating :
X arm beat frequency = 73.2 MHz; Y arm beat frequency = 26.6 MHz.

Attachment

1.the ALS in-loop noise in X and Y arms with IR off resonance and resonating.

2.the ALS in-loop noise in Y arm in each step from 1 to 6.(will follow soon)

Discussion

The Y arm ALS in-loop noise doesn't seem to be different in any of the configurations in step 1 to 6. This seems to mean that the ALS of the two arms are decoupled.

Actually we are not sure what changed from the last few days (when we were seeing some sort of coupling between the ALS of X and Y arm) except for YARM green PDH servo gain changed (see this entry),

Attachment 1: ALS_XY_inloop.pdf
ALS_XY_inloop.pdf
  9198   Thu Oct 3 13:54:25 2013 manasaUpdateGreen LockingALS out-of loop noise

We found the PDH servo gain for Y arm green was set at 2 (too low). The gain was set to 8.6 (based on earlier OLTF measurement elog 8817).

The ALS out-of loop noise was remeasured. We also measured the out-of loop noise of each arm while the other arm had no green (shutter closed). There doesn't seem to be any difference in the noise (between green and orange for Y arm and red and pink for the X arm) except that the noise in the X arm was slightly low for the same conditions (blue and red)  when measurement was repeated.

TRANSLATION by Jenne:  We first locked both X and Y for IR using the LSC, and X and Y for green using the analog PDH servos.  We measured the _PHASE_OUT_Hz calibrated error signals for both X and Y in this configuration - this gives us the out of loop noise for the ALS system, the Green and Blue traces in the plot.  We then closed the X end shutter, and measured the Y arm's error signal (to check to see if there is any noise contribution from the suspected X-Y cross beatnote).  Then, we closed the Y end shutter, relocked the Xarm on green's 00 mode, and measured the X arm's error signal.  We weren't sure why the Pink curve was smaller than the Blue curve below a few Hz, so we repeated the original measurement with both arms dichroic.  We then got the Red curve.  So, we should ignore the blue curve (although I still wonder why the noise changed in such a short time period - I don't think we did anything other than unlock and relock the cavity), and just see that the Green and Gold curves look similar to one another, and the Red and Pink curves look similar to one another.  This tells us that at least the out of loop noise is not affected by any X-Y cross beatnote.

Attachment 1: ALS_outloop.pdf
ALS_outloop.pdf
  9197   Thu Oct 3 10:29:03 2013 masayukiUpdateGreen LockingNew ALS autolocker flowchart

 

 [Manasa, Masayuki]

We made a new flowchart of ALS autolocker. We added the additional step to find the beat note frequency. We have to find a way to read the PSL temperature. By reading the PSL temperature we can decide the sweep range for the end green laser temperature with the curve which measured in previous measurement (in this entry)

We have three thresholds of error signal. One is the threshold for checking the arms are stabilized or not. It should be some hundreds count. Another threshold is to check that the suspensions are not kicked. This should be some thousands counts (in flow chart, it is 2K counts). The other is to check the optimal servo gain. If the servo gain is too high, the UGF is also too high and we will not have enough gain margin. The error signal start to oscillate at the UGF. We will check this oscillation and find the optimal gain. In flow chart this threshold is 1K counts.

Attachment 1: flowchart.pdf
flowchart.pdf
  9196   Thu Oct 3 09:49:49 2013 SteveUpdateSUSETMY oplev laser is sick

Quote:

As another proof that sometime is ill with ETMX Optical Lever:

We scanned the ETMX bias in PIT using ezcastep and saw that the OL response is very screwy. In the attached, you can see that the ETMX SUSPIT signal shows that the actual motion is good and linear. In fact, our sus diagonalization is extermely good and there's almost no signal in SUSYAW.

 ETMY oplev laser clearly showing a tail when it was projected up the sealing.

PS (10-4-2013): I checked the beam quality again as it was removed from the table: it had a good image at 3 meters

  9195   Thu Oct 3 09:01:06 2013 manasaUpdateGreen LockingALS high frequency noise

 As I was trying to solve the 2 arm ALS problem, I found the Y arm ALS not so stable AGAIN :( . I measured the in-loop noise of the X arm as ~400Hz/rtHz (60 picometers).

I went ahead and checked the out of loop noise of the ALS and found there is some high frequency noise creeping in above 20Hz for the Y arm ALS (blue curve). I checked the UGFs and phase margins of the phase tracker loops and found they were good (UGF above 1.4KHz and phase margins between 40 and 60 degrees).

So the suspect now is the PDH servo loop of both the arms which has to be checked.

Attached is the out-of loop noise plots of X and Y arm ALS.

Attachment 1: ALS_outloop.pdf
ALS_outloop.pdf
  9194   Thu Oct 3 08:57:00 2013 jamieUpdateComputer Scripts / Programspianosa can't find Jamie PPA

Quote:

Message on 'pianosa':

Failed to fetch http://ppa.launchpad.net/drgraefy/nds2-client/ubuntu/dists/lucid/main/binary-amd64/Packages.gz  404  Not Found

Sorry, that was an experiment to see if I could set up a general-use repository for the NDS packages.  I've removed it, and did an update/upgrade.

  9193   Thu Oct 3 02:51:26 2013 JenneUpdateLSCPRMI locked "forever", some ALS fiddling

First up for me this evening was getting the PRMI locked.

I used the IFO configure screen to lock the X and Y arms, then aligned them using the ASS scripts.  Then used the IFO config screen to restore the Michelson, and did some fine tune tweaking of the BS alignment by looking at the AS camera.  Then, I restored the PRMI from the IFO config screen, tweaked the PRM a little bit in yaw, and was able to get a lock using REFL 165 I&Q for ~25 minutes before I got bored and unlocked things.  I used the ASS for the PRM to align the PRM, then turned off the ASS.  POP110 and POP22 both drifted down, but by a small amount, and at the end (when I turned the ASS back on for PRM), they picked back up to about their original levels.

PRMI_lock_ASSatBeginningAndEnd.png

(Note to self: to get it to print both plots, chose custom paper size, make it 14.5 by 11.  Don't ask why, just do it, because it works.  Also, in PNG device properties, increase the compression to 9.)

 

After I played with the PRMI, I started looking at the ALS system. 

I had both arms locked on IR using the regular LSC system (so POX and POY for the error signals).  Then I opened up the green shutters, and got both arms locked on green (so the green lasers were just following the arms...no digital ALS business).  I went out to the PSL table and tweaked up the alignment of the green beams (didn't need much at all, just an itsy bitsy bit in yaw, mostly).  I saw a very strong peak for the Yarm vs. PSL (around -19dBm), and there was a harmonic of that beat.  Opening and closing the Xarm green shutter had no effect on these peaks, so there wasn't any kind of X-Y cross beat sneaking around that I could see.  That's really as far as I got - I think (but haven't checked) that Manasa may have removed the power splitter / combiner, so that the RF analyzer is only looking at the Y beat PD (she mentioned earlier today that she was going to give that a try to narrow things down). 

After that, Rana and I went back to the PRMI for some noise stuff, and worked on the PMC.  See those separate elogs for info on those activites.

  9192   Thu Oct 3 02:46:58 2013 ranaMetaphysicsPEMUSGS Furlough

ObamaCare_GOP.png

  9191   Thu Oct 3 02:43:34 2013 rana, jenneSummaryLSCPRMI: comparison of 1f and 3f signals w/ calibration

The attached plot shows the spectra of all the REFL signals with the PRMI SB lock.

We excited the ITMY_LSC with 3000 counts. We used the Masayuki calibration of ITMY (5 nm / count * (1/f^2)) to estimate this peak in the REFL spectra.

To correctly scale the REFL spectra we account for the fact that the DTT BW was "0.187 Hz" and we turn off the "Bin" radio box before measuring the peak height with the cursor.

Since the ITMY motion is 3000 * 5e-9 / (580.1 Hz)^2 = 44.6 pm_peak, we want the DTT spectrum of the REFL spectra to report that too.

i.e. to convert from peak height to meters_peak, we use this formula:

meters_peak = peak_height * sqrt(BW) * sqrt(2)

I *think* that since the line shows up in multiple bins of the PSD, we should probably integrate a ~0.5 Hz band around the peak, but not sure. Need to check calibration by examining the time series, but this is pretty close.

Mystery: why are the REFL_I 3f signals nearly as good in SNR as the 1f signals? The modelling shows that the optical gain should be ~30-100x less. Can it be that our 1f electronics are that bad?

Bonus: notice how we have cleverly used the comb of bounce frequencies around the calibration line to determine that REFL11 is clipping!

Attachment 1: REFL_signals_CalLinesLinedUp.pdf
REFL_signals_CalLinesLinedUp.pdf
  9190   Thu Oct 3 01:24:31 2013 Jenne, RanaUpdateIOOPMC

The PMC transmission was around 0.78 all day, rather than the usual 0.83ish.  Rana went out to the PSL table and fixed up the PMC alignment.  This should not need to be done very often, so things to check before touching the alignment are FSS / PMC settings (digital stuff).  Make sure that the PC RMS (on the FSS screen) is low (at least below 2, preferably below 1), and that the FSS Fast monitor is near 5ish (not near 0 or 10).  

This is a capture of PMC REFL's camera after Rana was finished. If it doesn't look this good when you finish then you are not done. Never do PMC alignment without looking at the PMC REFL camera.

PMCR_1064822387.bmp

The attached trend shows 80 days of PMC REFL and TRANS. The bad alignment stuff started on Sep 21-24 time period. You know who you are.

Attachment 2: pmc80.png
pmc80.png
  9189   Thu Oct 3 01:18:57 2013 KojiHowToLSCsteps to full IFO

I vote on PRMI+1arm -> PRFPMI

  9188   Thu Oct 3 01:06:48 2013 rana, jenneUpdateSUSoplev XY-plots reflect new calibration

As another proof that sometime is ill with ETMX Optical Lever:

We scanned the ETMX bias in PIT using ezcastep and saw that the OL response is very screwy. In the attached, you can see that the ETMX SUSPIT signal shows that the actual motion is good and linear. In fact, our sus diagonalization is extermely good and there's almost no signal in SUSYAW.

Attachment 1: etmx_ol.png
etmx_ol.png
  9187   Thu Oct 3 00:01:59 2013 rana, jenneHowToLSCsteps to full IFO

In moving now to full IFO locking, there are a number of sub-states to diagnose:

  1. PRMI + 1 arm
  • Measure sensing matrix as arm is scanned into resonance. Compare time series of sensing matrix elements with New LoopTickle simulation. But first, we need more than 1 LOCKIN screen in the LSC! That will allow us to measure all of the elements of front_matrix.jpg simulataneously.
  • Measure 3f PRMI noise spectra as a function of arm position. Look for trouble.
  1. DRMI + 1 arm
  • Same as PRMI above.
  • Want to find why this is unstable sometimes. Make stable for t > 10 minutes.
  • Maybe add some QPD->ASC for SRC angular control, but how? Will this still work after the arms are resonant or will it be swamped by carrier contrast defect? Will Berlusconi ruin all of the Italian gelateria? Only time can tell...
  1. FPMI (non optically recombined) for ALS diagnosis
  2. PRFPMI (iLIGO configuration)
  • this ought to be easier than DRFPMI
  • will let us tell if our ALS is good enough to handle the coupled cavity pole
  1. DRFPMI (aLIGO style)

Which to do first and in what order?

  9186   Wed Oct 2 23:21:54 2013 ranaUpdateComputer Scripts / Programspianosa can't find Jamie PPA

Message on 'pianosa':

Failed to fetch http://ppa.launchpad.net/drgraefy/nds2-client/ubuntu/dists/lucid/main/binary-amd64/Packages.gz  404  Not Found

  9185   Wed Oct 2 14:25:39 2013 SteveUpdateVACvertex crane is still out of order

 

The smoke alarms were turned off and surrounding areas were covered with plastic.

The folding I-beam was ground down to be in level with the main beam.

Load bearing cable moved into correct position. New folding spring installed.

Crane calibration was done at 500 lbs at the end of the fully extended jib.

Than we realized that the rotating wheel limit switch stopped working.

This means that the crane is still out of order.

 

Attachment 1: grindingIbeam.jpg
grindingIbeam.jpg
Attachment 2: cableMoved.jpg
cableMoved.jpg
Attachment 3: 500lbs@end.jpg
500lbs@end.jpg
  9184   Tue Oct 1 19:42:19 2013 ranaSummaryCDSmegatron upgrade

Max and I started upgrading megatron to Ubuntu 12.NN today. We were having some troubles with getting latest python code to run to support the Summary pages stuff.

Its also a nice test to see what CDS tools fail on there, before we upgrade the workstations to Ubuntu 12.

Since its Linux, none of the usual upgrading commands worked, but after an hour or so of reading forums we were able to delete some packages and all the 3rd party packages and get the upgrade to go ahead. We'll have to re-install the LSC, GDS, LAL repos to get it back into shape and get NDS2 working. The upgrade is running in a 'screen' command on there.


Wed Oct 02 14:50:16 2013 

Update #1: The upgrade asks a couple dozen questions so it doesn't proceed by itself. I've been checking in to the 'screen' every couple hours to type in 'Yes' to let it keep going.


Update #2: It finished a few hours ago:

controls@megatron:~ 0$ uname -a
Linux megatron 3.2.0-54-generic #82-Ubuntu SMP Tue Sep 10 20:08:42 UTC 2013 x86_64 x86_64 x86_64 GNU/Linux
controls@megatron:~ 0$ date
Wed Oct  2 18:33:41 PDT 2013

  9183   Tue Oct 1 17:14:53 2013 masayukiUpdateGreen LockingALS servo filters modified

 

 [Manasa, Masayuki]

[revised at 10/1 pm 5:00]

As we mentioned in previous entry (elog#9171), the phase margin of ALS control was at most 20 degree. We modified the filter of C1ALS_XARM and C1ALS_YARM. The OLTF is in attachment1. Now the phase margins of both arms are more than 35 degree. I modified the FM5 filters of both servo.

FM5 filter is the filter for the phase compensation. It had the one pole at 1000 Hz and one zero at 1Hz. As you can see in attachment2, it start to lose the phase at 50 Hz. But the UGF of our ALS control loop is higher than 100 Hz, so I changed the pole from 1 kHz to 3 kHz in order to get more phase margin at UGF. The new servo have 10dB larger gain than previous filter at higer than 1kHz, but the control loop do nothing in that region, so it's no problem.

We have phase lag between 2 arms. I used same filters for both arms, so I'm wondering where these phase lag came from.

 

Attachment 1: OLTF.pdf
OLTF.pdf
Attachment 2: filter_change.pdf
filter_change.pdf
  9182   Tue Oct 1 14:12:22 2013 ranaSummaryCDSsvndumpfilter on linux1 makes NFS slow

 Yesterday and this morning's slow NFS disk access was caused by 'svndumpfilter' being run at linux1 to carve out the Noise Budget directory. It is being moved to another server; I think the disk access is back to normal speed now.

  9181   Tue Oct 1 11:24:02 2013 ranaUpdatePEMparticle counts

 Probably more important is to establish quantitatively how particle counts affects the lock acquisition or noise in the interferometer.

We don't want to adopt a "Sky is Falling" mentality as was done previously in LIGO when people were trying to outlaw burritos and perfume.

Dust monitor counts and human noses do not correlate well with the interferometer's nose.

pepelepew.jpg

When the vacuum system is closed, we might check to see if particle counts correlate with losses in the PMC or excess scatter on the ISC tables. If not, we should move on to other concerns.

  9180   Tue Oct 1 10:06:27 2013 SteveUpdatePEMparticle counts

How can we improve the cleanliness of the 40m IFO room 104 ? May be Mario Batali crocs if you chief chef of the lab likes it. We can do more effective things.

Atm 1 is showing the lab air quality dependence on outside air  measured at the top of the IOO chamber. This made me to measure the   differential  pressure between lab and out side.

Properly sized air conditioner is over pressurizing it's room by 0.05 [" Water] like clean assembly room.

The 40m at Vertex location is barely getting pressurized to 0.01 " W

The control room is OK with 0.04 " W but it's air quality is very bad! It is 10x worse at 0.5 micron than IFO room.

Every entry from well pressurized control room into barely pressurized  IFO  pumps dirty air to our "clean room."

This door should be closed.

The drill- room entry is ideal because it is using the same air conditioner as the  main lab, therefore it has cleaner air.

 

Things to do: seal holes of CES walls, seal paint wall at south end so it will shed less, replace gas kits on doors.

I have plugged the cable tray space entering the control room above Rosalba.

Attachment 1: 3xJuly4Cal.png
3xJuly4Cal.png
Attachment 2: whichOne.jpg
whichOne.jpg
  9179   Tue Oct 1 09:51:10 2013 ranaUpdateGreen LockingALS autolocker flowchart

  I think we can use the IMC autolocker to start with getting this started. Once Jamie fixes the NDSSERVER environment variable bug, we should be able to use his more slick automation code to make it auto lock.

  9178   Mon Sep 30 23:56:19 2013 manasaUpdateGreen LockingALS autolocker flowchart

[Masayuki, Manasa] 

Flowchart for ALS autolocker. The error signal thresholds will be decided by trial and error.

 ALSautolocker.png

  9176   Mon Sep 30 17:55:45 2013 manasaUpdateGreen LockingX and Y arm transmission needs to be decoupled

[Masayuki, Manasa]

Problem
We wanted to lock both the arms using ALS and get IR to resonate while arms are held using ALS. The X arm was locked using ALS and offsetter2 was used to scan the arm and find IR resonance. The Y arm was locked using ALS. But as the Y arm was brought closer to IR resonance, the X arm ALS loses lock. (attachment 1)

Discussion
We believe that this comes from the X and Y transmission not being well separated at the PSL table. The PBS is not sufficient to decouple them (A strong beatnote ~35dB between the X and the Y arm green lasers can be seen on the spectrum analyzer).

Solution
Decouple the X and Y arm transmitted beams at the PSL table. I am trying to find a wedged mirror/window that can separate the 2 beams at the PSL table before the beat PD (sadly the laseroptik HR532nm optics have no wedge)

 

Attachment 1: scan2.png
scan2.png
  9175   Mon Sep 30 13:02:51 2013 Masayuki,ManasaUpdateIOOPMC and MC alignment

[Manasa, Masayuki]

The MC lost lock around 8+hrs ago. The transmission from PMC was 0.77 this morning, so we aligned the PSL to the PMC using the two steering mirrors before the PMC. We brought the PMC transmission to 0.841. We also aligned the MC, and the MC transmission reflection now is 0.59.

  9174   Mon Sep 30 11:33:15 2013 ranaUpdateLSCLSC calibration screen

Quote:

  I fixed the XARM and YARM real time calibration servo.

I also change the C1CAL_MICH_A servo. Now the actuator response and the suspension TF are combined together and that filter name is BS_act. C1CAL_XARM_A and C1CAL_YARM_A have same kind of filters, ETMX_act and ETMY_act.

There are AI filter in each A servo and inv_AA, inv_DAA filters in CINV servo, but it's doesn't work correctly yet.

 These aren't servos. What he means is that he's changed some filters in the real time calibration screens so as to make the actuation and sensing parts more accurate, but the inversion of the AA filters is not accurate yet.

  9173   Mon Sep 30 08:53:22 2013 SteveUpdateVACvertex crane repair schedule has changed

Quote:

Quote:

 The folding beam removed as shown. Two man supporting it while I hammering it out. Pin was dry and it gulled into supporting hinges.

The rotating hinge will be machined and bushing will be added with Zerk fitting or similar. This will allow lubrication in the future.

 see elog #9111

Atm1,  The folding arm is back on with 0.1" misalignment at no load in the trolly's way. The other side of the I beam is 0.02" higher than the main beam.. New bushing and pin were greased up  with Krytox before installation.

The axial Zerk 1/8" pipe in the pin upper end can not take any fitting. There is no room. It is taped off.

This gap comes down to ~ 1/16" at fully extended arm with 225 lbs load at the end of it.

The present plan is to grind down the  the misalignment of 0.1"  for a slow-loaded trolly.

Steve Baker of Konecranes will be back to grind down this ridge and load test at 500 lbs on Tuesday,  OCT  1, 2013

 

 

 

 Konecranes rescheduled the completion of the Vertex crane to Wednesday, Oct. 2

  9172   Fri Sep 27 21:01:50 2013 MasayukiUpdateLSCLSC calibration screen

 

 I fixed the XARM and YARM real time calibration servo.

I also change the C1CAL_MICH_A servo. Now the actuator response and the suspension TF are combined together and that filter name is BS_act. C1CAL_XARM_A and C1CAL_YARM_A have same kind of filters, ETMX_act and ETMY_act.

There are AI filter in each A servo and inv_AA, inv_DAA filters in CINV servo, but it's doesn't work correctly yet.

  9171   Fri Sep 27 20:28:10 2013 manasaUpdateGreen LockingALS servo

[Masayuki, Manasa]

I. ALS servo loops
After fixing things with the phase tracking loop, we checked if things were good with the ALS servo loops.
We measured the OLTF of the X and Y arm ALS servo loops. In both cases the phase margin was ~20 degrees. There was no room to set enough phase margin. So we looked at the servo filters. We tried to modify the filters so that we could bring enough phase margin, but could not get at it. So we put back the old filters as they were.

 attachment1: OLTF of the ALS XARM and YARM control loops

attachment2: Current phase budget. FM4 and FM10 are the boost filters.

II. ALS in-loop noise
Also, I found that the overall noise of the ALS servo has gone up by about two orders of magnitude (in Hz/rtHz) over the whole range of frequencies for both the arms from the last time the measurements were made. I suspect this could be from some change in the calibration factor. Did anybody touch things around that could have caused this? Or can somebody recollect any changes that I made in the past which might have affected the calibration? Anyways, I will do the calibration again.

 

 

Attachment 1: OLTF.pdf
OLTF.pdf
Attachment 2: phase_badget_xarm_ALS.pdf
phase_badget_xarm_ALS.pdf
  9170   Fri Sep 27 16:02:23 2013 manasaUpdateGreen LockingY arm ALS phase tracking loop gain changed

[Masayuki, Manasa]

While trying to lock the arms using ALS we found that the locks were not very stable and the in-loop noise was higher than seen before.

I looked into things and checked the out-of loop noise for ALS and found that the Y arm ALS noise (rms) was higher than the X arm.

To troubleshoot, I measured the OLTF of the phase tracking loop. While X arm was healthy, things weren't looking good for the Y arm. Sadly, the Y phase tracking loop gain was set too high with a phase margin of -2 degrees. We brought down the gain from 300 to 150 and set the phase margin close to ~55 degrees.

X arm Phase tracker loop:
UGF = 1.8 K Hz
Phase margin = 50 degrees

Y arm Phase tracker loop:
UGF = 1.6 KHz
Phase margin = 55 degrees

Attachment 1: outofloop.pdf
outofloop.pdf
Attachment 2: PTX_OLTF.pdf
PTX_OLTF.pdf
Attachment 3: YPT_OLTF_after1.pdf
YPT_OLTF_after1.pdf
Attachment 4: YPT_OLTF_before.pdf
YPT_OLTF_before.pdf
  9169   Fri Sep 27 13:55:11 2013 MasayukiUpdateLSCLSC calibration screen

  I added the DAQ channel to all output of calibration servo. The name of channels are C1CAL_(plant name)_W_OUT_DQ.

I recompiled and restarted the model. Also I committed the changes to the svn of the calibration model.

  9168   Fri Sep 27 00:48:53 2013 MasayukiUpdateLSCFPMI noise caused by ARM locking

Quote:

Hidden in Nakano-kun's previous entries was that the phase margin of the X-Arm was only 9 degrees!! This extremely close to instability and makes for huge gain peaking. The feedback loop is increasing noise above 100 Hz rather than suppress. After some tweaks of the LSC filters we got a much more stable loop/.

So we today started to examine the sources of phase lag in the arm cavity sweeps. There were a few unfortunate choices in the XARM LSC filter bank which we tuned to get less delay.

Then I wrote a bunch of detail about how that worked, but the ELOG ate my entry because it couldn't handle converting my error signal noise plot into a thumbnail. Then it crashed and I restarted it. We also have now propagated the changes to the Y arm by copy/paste the filters and the result there is pretty much the same: low phase margin is now 38 deg phase margin. Noise is less bad.

 

[Rana, Masayuki

 I made the plot of the phase of the digital filters which Rana change and also  of the AA, AI, DAA, DAI filters. Now the biggest phase delay come from the timedelay of the digital system.

phase_badget.png

The UGF is around 150 Hz at that frequency the time delay has biggest phase  delay. Second one is the FM9 filter (this filter is BOOST filter). Then we have the AA filter, AI filter and so on, but these delay is roughly 5 degree.

As I said in previous entry, the time delay of the XARM control is roughly 300 usec, and we have 120 usec even only in C1SUS. Also between the C!SUS and C1LSC we have another 120 usec time delay. We want to increase the UGF to 300 Hz but because of the time delay of the digital system we cannot increase. So we should fix this problem.

 

After changing these filters, the FPMI noise is become better at high frequency. Before we have peak around the 100 Hz (because of 8 degree phase margin...), but they are gone. i attached the noise spectrum. This plot is measured by the real time calibration output. But even then, you can see the extra noise around 100 Hz in FPMI conpare to only MICH.

 FPMI_noises.png

 

 

  9167   Thu Sep 26 23:02:40 2013 ranaUpdateLSCFPMI noise caused by ARM locking

Hidden in Nakano-kun's previous entries was that the phase margin of the X-Arm was only 9 degrees!! This extremely close to instability and makes for huge gain peaking. The feedback loop is increasing noise above 100 Hz rather than suppress. After some tweaks of the LSC filters we got a much more stable loop/.

So we today started to examine the sources of phase lag in the arm cavity sweeps. There were a few unfortunate choices in the XARM LSC filter bank which we tuned to get less delay.

Then I wrote a bunch of detail about how that worked, but the ELOG ate my entry because it couldn't handle converting my error signal noise plot into a thumbnail. Then it crashed and I restarted it. We also have now propagated the changes to the Y arm by copy/paste the filters and the result there is pretty much the same: low phase margin is now 38 deg phase margin. Noise is less bad.

 

Attachment 1: Xarm_sweep_130926.pdf
Xarm_sweep_130926.pdf
Attachment 2: lsc.pdf
lsc.pdf
Attachment 3: err.png
err.png
  9166   Thu Sep 26 21:55:08 2013 ranaUpdateSUSProblems with ETMY Optical Lever

Not so fast! We need to plan ahead of time so that we don't have to repeat this ETMY layout another dozen times. Please don't make any changes yet to the OL layout.

Its not enough to change the optics if we don't retune the loop. Please do buy a couple of JDSU (and then we need to measure their intensity noise as you did before) and the 633 nm optics for the mode matching and then we can plan about the layout.

  9165   Thu Sep 26 11:00:51 2013 SteveUpdateSUSProblems with ETMY Optical Lever

We are out of JDSU-Uniphase 1103P heads. I'm ordering some right now. I'm planning to make some corrections on Rana's list tomorrow morning at ETMY.

  9164   Thu Sep 26 10:55:29 2013 SteveUpdateGeneralY-arm floor space for test

Liyuon will set up a  ~5 mW He/Ne laser for  waist measurement for LIGO oplev telescope.

This  will be between the beam tube and the CES wall. He will do his tests in the morning.

  9163   Thu Sep 26 01:49:28 2013 MasayukiSummaryGreen LockingFPMI noise caused by ARM locking

 

 Measurement with FPMI


i)By locking the FPMI with AS55Q and arms using POX,POY we measured  the OLTF on AS55Q, the response from BS actuation to error signal on AS55Q  for H_mich. The fitted,  measured OLTF and the residual function is in attachment1. I fitted two parameters and they are time-delay and the gain. The time delay is -275 usec. The time delay in three different control are almost same. The response from BS to AS55Q is in attachment 2.


With these two measuremets, I calclated the H_mich in FPMI. This H_mich should be different from simple MI because the cavity  refrectivity is different from the front mirror. Acrually it changed and the value was
Hmich = 4.4026e7

ii) I excited the ETMX and ETMY and measure the response from actuation to the error signal of MICH on AS55Q. The response is in attachment 3 and 4. from these result I calculated the H_L-l by using the formula as I mentioned. The value was
H_Lx-l = 175.7650 (XARM)
H_Ly-l = 169.8451 (YARM)

iii) I measured the error signal of MICH and XARM and YARM and with measured H_L-l, I estimated the FPMI noise caused by ARM locking. You can see in the higher frequency region than 10 Hz is dominated by noise caused by ARM control in-loop noises. 150 Hz and 220Hz are the UGF of each arms, so the two peaks are caused by arm control. You can see the small difference between FPMI noise and  noise from arms. There are two possibilities, one is that these measurement is not same time measurement so they should have small difference. and  other possibility is the error of the caliculation. But I think it doesn't look so bad estimation.

 

Next step

We will do same measurement with lock the arms the ALS system on tomorrow. Then we will check the PDH servo or other noise source and investigate the ALS system

Attachment 1: MICH_OLTF.pdf
MICH_OLTF.pdf
Attachment 2: BS-RS55Q.png
BS-RS55Q.png
Attachment 3: ETMX-RS55.png
ETMX-RS55.png
Attachment 4: ETMY-RS55.png
ETMY-RS55.png
Attachment 5: plot.pdf
plot.pdf
ELOG V3.1.3-